I need help converting some things from moles to grams, and grams to mL. However, some of these calculations need to be done using mass ratio...

Convert 0.0537 mole of 30% hydrogen peroxide. **mass ration** - 30% by mass means that you have 30g of H2O2 per 100g of solution. Calculate the mass (g) and then convert to mL.

A solution of Cyclohexene (88.8% by mass) and the phase transfer catalyst Aliquat 336 (11.2% by mass).

0.0122 mole of cyclohexene . Calculate the number of grams of cyclohexene, then the mass of the cyclohexene/Aliquat solution using the data given above.

Thank you :)
Answer #1

1) The molar mass of hydrogen peroxide is 34.0147 g/mol, implying that 34.0147 g of the compound will make up 1 mol of it. So, 0.0537 mol of the compound will be 0.0537*34.0147 = 1.8265 g.

Next, since the solution is only 30% hydrogen peroxide by mass, (3/100) of the total solution is 1.8265g. Therefore, the total volume of the solution will be, by a simple unitary method 1.8265*(100/3) = 60.8833 g.

Finally, the density of a 30% hydrogen peroxide solution is 1.11 g/mL. So, 60.8833 g of the solution will occupy 60.8833/1.1154.8498 mL.

2) The solution of cyclohexene/Aliquat 336 given, contains 88.8% by mass cyclohexene, which is 0.0122 mol, the mass of cyclohexene is 0.0122*82.143 = 1.0021 g. Since 1.0021 g is 88.8% of the solution, the total mass of the solution will be 1.0021*(100/88.8) = 1.1285 g. 11.2% by mass of this solution is Aliquat 336. So, that will be (11.2/100)*1.1285 = 0.1264 g of Aliquat 336.

Therefore, summarizing, the total mass of the solution will be 1.1285 g with the individual masses of cyclohexene and Aliquat 336 being 1.0021 g and 0.1264 g respectively.